OpenStudy (nsh4267):

A construction firm is trying to decide which of two models of a crane to purchase. Model A costs $100,000 and requires $8000 per year to maintain. Model B has an initial cost of $80,000 and a maintenance cost of $11,000 per year. For how many years must model A be used before it becomes more economical than B?

OpenStudy (anonymous):

CostA = Years*8000 + 100000 CostB = Years*11000+ 80000 Now we solve CostA = CostB 8000Y + 100000 = 11000Y + 80000 http://www.wolframalpha.com/input/?i=8000Y+%2B+100000+%3D+11000Y+%2B+80000

OpenStudy (anonymous):

Graph included in the link

OpenStudy (nsh4267):

THANK YOU!!!